Suppose that the distribution of weights of Philippine hamsters is normally distributed with a mean of 65.5 g and standard deviation of 11.6 g. If there are 1000 weights in this population, how many of them are greater than or equal to 79 g?


1
Expert's answer
2021-05-12T01:53:06-0400

"P(X>\frac{79-65.5}{11.6})=P(Z>1.16)=1-P(Z<1.16)=1−0.8770=0.1230."

Among 1000 weights there are "1000*0.1230=123" greater than 79 g.